calculate the longest length of steel wire that can hang vertically with out breaking .breaking stress for steel=7.982*10^8N/m^2 and density of steel(d)=8.1*10^3kg/m^3.

Grade:11

2 Answers

Nirmal Singh.
10 years ago
Breaking Stress = 7.982x102 N/m2
Mass m= Volume x density = V x d
for no breaking F = mg = V x d x g = A x L x d x g
Lets consider Longest length is L and Area is A
Now Stress = F/A
Stress = A x L x d x g / A
Stress = L x d x g
7.982x102 = L x 8.1×10³ x 9.8
L = 0.01006 m
L = 10.06 mm
Thanks & regards,
Nirmal Singh
Askiitians faculty
Nirmal Singh.
10 years ago
Breaking Stress = 7.982*10^8N/m^2,
density of steel(d)=8.1*10^3kg/m^3
lets consider Longest length is L and mass is m area is A
Stress = F / A
For nt breaking F = mg = V x d x g = A x L x d x g
Stress = A x L x d x g / A
Stress = L x d x g
7.982*10^8 = L x 8.1*10^3 x 9.8
L = 10 m
Thanks & regards,
Nirnal Singh
Askiitians Faculty